#1 The market ties are competitive and demand is given by P=210-Q. If supply is given by P=2Q, what will be the price in this market and what quantity?

#2  Suppose that the administration requires all of its faculty and students to wear new ties every day. Draw a supply and demand graph of this situation indicating what will happen to the market price and quantity for the ties. Why does the price change in this market?

#3 When P=5, quantity is Q=100 When P=10, quantity is Q=55   What is the elasticity of demand? Show your work. What are the factors that determine the elasticity of demand? Just list these factors
Answer #1

1 [ Demand p = 210 - Q Supply, p=20 Im equilibrium, Demand = Supply 210 - Q = 28 2 Q = 70 P = 140 ᏁᎢ , ᏗᎩ , , Ꭻ , ᎰᎢ ᎬᏔ ) 11-- This right ward shift of demand come will increase both price and quantity of tie in the equilibrium. Price will change be